OpenStudy (anonymous):

integral from 0 to 1 of (t(e^(-t^2))i+tln((t^2)+1)j) dt @ganeshie8 hello? ._.

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \left( te ^{-t ^{2}} i+tln \left( t ^{2} +1 \right)j\right) dt\]

OpenStudy (dan815):

u sub on the i component and int by parts on the j

OpenStudy (anonymous):

so separate the integrals. then u sub for i and IBP for j. then just combine into a vector, right?

OpenStudy (dan815):

yea derivative of a vector is a vector and integral of a vector is a vector

OpenStudy (anonymous):

@ganeshie8 I understand what he said to do but im just having computational problems :(

ganeshie8 (ganeshie8):

heyy

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{1} \left( te ^{-t ^{2}} i+t\ln \left( t ^{2} +1 \right)j\right) dt\] is same as \[\left(\int\limits_{0}^{1} te ^{-t ^{2}} dt~\right)i+\left(\int\limits_{0}^{1} t\ln ( t ^{2} +1)~dt\right) j\]

ganeshie8 (ganeshie8):

evaluate each integral and plugin ?